OpenStudy (chrisplusian):

HELP PLEASE!!!!! Find the vectors T, N, and B at the given point... r(t) = 7cost, 7sint, 7lncost (7,0,0)

OpenStudy (chrisplusian):

For the unit tangent vector I got \[T(t)= <sintcost, \cos^2 t,sint>\]

OpenStudy (chrisplusian):

Then evaluating for T(0) I got <0,1,0> and webassign said that answer was correct. I evaluated T(0) because I set the i-th component equal to the x value of the point. Then I set the j-th component equal to the y value of the point and both indicate that the value of t=0-degrees.

OpenStudy (chrisplusian):

For the derivative I got \[T'(t)= <\cos^2 t-\sin^2 t/-2costsint, cost>\]

OpenStudy (chrisplusian):

If I vealuate T'(0)= <1,0,1> so I use that to find the magnitude which is just the sqaure root of two. When I divide the vector through by that I get the unit normal vetoer "V(t)"\[V'(t)= < \frac{ 1 }{ \sqrt{2} },0, \frac{ 1 }{ \sqrt{2} } >\]

OpenStudy (chrisplusian):

My online homework liked the first answer for T(t) but not the second for V(t)...... and that last posting should not have been V'(t) but V(t)

OpenStudy (loser66):

@Psymon

OpenStudy (loser66):

what is V'(t)? unit vector of T?

OpenStudy (chrisplusian):

no it is the unit normal vector...... found by taking the derivative of the unit tangent vector and dividing it by the magnitude of the derivative of the unit tangent vector

OpenStudy (loser66):

so, you should use N as its symbol

OpenStudy (chrisplusian):

I meant to but I accidentally put that and then I retyped it twice making errors on both tries and finally I got everything in the equation accept the N and I write v' instead but yes I agree it should be N

OpenStudy (loser66):

I think you make mistake at N, you cannot take T' =<1,0,1> and |T'| = 1/sqrt 2..... because it's not that . you should do something like \[N=\frac{T'}{|T'|}=\frac{<cos 2t, sin2t, cost>}{\sqrt{cos^2(2t)+ sin^2(2t)+ cos^2t}}\]

OpenStudy (loser66):

and then plug t to calculate N.

OpenStudy (chrisplusian):

I actually agree. But remember we found t=0 or 180 degrees. If you plug that in and calculate it I seem to get that same result. Do you?

OpenStudy (loser66):

oh , yea, you are right

OpenStudy (chrisplusian):

However the homework assignment disagrees with you on that point

OpenStudy (chrisplusian):

It says that this is not the answer

OpenStudy (chrisplusian):

@theEric

OpenStudy (chrisplusian):

Any Idea?

OpenStudy (theeric):

No, sorry... I tried a little bit, but I got a different \(\vec T(t)\) from yours, and you said yours was correct. \(\vec r(t) = \left<7\cos t,\ 7\sin t,\ 7\ln(\cos t)\right> \\\implies \vec r'(t)=\left< -7\sin t,\ 7\cos t,\ 7\tan t\right> \\\implies \dfrac{\vec r'(t)}{\left|\left|\vec r'(t)\right|\right|}=\vec T(t)=\dfrac{\left< -7\sin t,\ 7\cos t,\ 7\tan t\right>}{\sqrt{49\sin^2 t+\ 49\cos^2 t+\ 49\tan^2 t}} \\~\\\qquad\qquad=\dfrac{\left< -7\sin t,\ 7\cos t,\ 7\tan t\right>}{7\sqrt{1+\tan t}}=\dfrac{\left< -\sin t,\ \cos t,\ \tan t\right>}{\sqrt{1+\tan t}} \)

OpenStudy (theeric):

Good luck!

OpenStudy (chrisplusian):

@theEric the only thing wrong with what you put in was that you started with tan^2 in the denominator then after you factored the 49 out you only wrote tan when it should have been tan^2 my paper looks identical to yours but with tan^2

OpenStudy (chrisplusian):

The 1+tan^2 would just be sec^2 and that comes out of the radical as just 7sec

OpenStudy (chrisplusian):

@ybarrap ?

OpenStudy (ybarrap):

yes, I agree @chrisplusian that it should be just 7sec t after taking square root in denom, but otherwise @theEric has the right solution for the unit tangent

OpenStudy (chrisplusian):

Ok so when I simplify that I get <-sintcost, cos^2(t), sint> do you agree?

OpenStudy (ybarrap):

yes!

OpenStudy (chrisplusian):

Ok so then the unit normal vector is the derivative of the unit tangent vector over the magnitude of the derivative of the unite tangent vector right?

OpenStudy (ybarrap):

yes

OpenStudy (ybarrap):

And you can check by taking dot product, should be zero

OpenStudy (theeric):

I see my mistake, thank you both! @chrisplusian yep!

OpenStudy (chrisplusian):

For the derivative of the unit tangent vector I got \[<\sin^2 t -\cos^2 t,-2sintcost, cost >\]

OpenStudy (chrisplusian):

do you agree with that?

OpenStudy (ybarrap):

yes, I get : <-cos(2 t), -2 sin(t) cos(t), cos(t)>

OpenStudy (theeric):

\( T(t)= \left<\sin t\cos t,\ \cos^2 t,\ \sin t\right> \) Looks good to me, too!

OpenStudy (chrisplusian):

Ok @ybarrap do you agree that this should be evaluated at t=0? from the original question?

OpenStudy (ybarrap):

yes

OpenStudy (ybarrap):

because cos(0)=1, sin(0)=0

OpenStudy (theeric):

\( \left<-cos(2 t), -2 sin(t) cos(t), cos(t)\right> \) @chrisplusian You do as we have, finding all of the vectors in terms of \(t\). Then you let \(t=0\). That looks good! :)

OpenStudy (chrisplusian):

Ok at this point can't I evaluate this and get T'(t)= < -1,0,1> ??

OpenStudy (chrisplusian):

excuse me T'(0)= <-1,0,1>

OpenStudy (chrisplusian):

@ybarrap ?

OpenStudy (ybarrap):

Don't you need to divide N by mag of N to make it a unit vector?

OpenStudy (chrisplusian):

I thought that\[N(t)=\frac{ T'(t) }{ \left| T'(t) \right| }\]

OpenStudy (theeric):

There might be a difference in ways of doing it. I do as @chrisplusian does, where \(\vec N\) is always a unit vector.

OpenStudy (chrisplusian):

Then if I evaluate T'(0) I come up with <-1,0,1> and the magnitude of the derivative of the unit tangent vector can be found by \[\sqrt{(-1)^2 +(0)^2 +1^2}=\sqrt{2}\]

OpenStudy (theeric):

I think you can do \(N(t)=\dfrac{ T'(0) }{ \left| T'(0) \right| }\) also.

OpenStudy (theeric):

I agree with this, @chrisplusian .

OpenStudy (ybarrap):

I agree as well.

OpenStudy (theeric):

:D

OpenStudy (chrisplusian):

So for the unit normal vector I came up with \[N(t)=<\frac{ -1 }{ \sqrt{2} },0,\frac{ 1 }{ \sqrt{2} }>\]

OpenStudy (theeric):

Looks good to me. :) What about the fun part, \(\vec B=\vec T\times\vec N\)

OpenStudy (ybarrap):

oh, yeah!

OpenStudy (ybarrap):

should be a snap now!

OpenStudy (chrisplusian):

Well the probem is that I have to put the answers in for each part before I can proceed to the next and web assign (an online homework) says that we are all three wrong and that is not the unit normal vector

OpenStudy (ybarrap):

I did the dot product between the two, and it is zero.

OpenStudy (chrisplusian):

@ybarrap will you explain what you are doing the dot product of?

OpenStudy (ybarrap):

T and N should be perpendicular

OpenStudy (chrisplusian):

the unit tangent vector and the original vector? And then the unit tangent vector and the unit normal vector?

OpenStudy (ybarrap):

the unit tangent and the unit normal should be perpendicular and their dot products must be zero. I checked, it is zero.

OpenStudy (theeric):

Huh!

OpenStudy (chrisplusian):

What about the unit tangent vector and the original vector? The should not be perpendicular should they?

OpenStudy (theeric):

Is it a formatting issue? Online questions are not smart..

OpenStudy (chrisplusian):

Smart enough to count as a grade :)

OpenStudy (ybarrap):

The original vector is neither tangent nor normal, generally.

OpenStudy (theeric):

It can vary. They're not strictly related, like \(\vec T\) and \(\vec N\). The \(\vec T\), \(\vec N\), and \(\vec B\) are all orthogonal. They make up a little coordiante system! :) http://www.math.tamu.edu/~tom.kiffe/calc3/tnb/tnb.html

OpenStudy (ybarrap):

The unit normal can be pointing in the wrong direction. Remember that the unit normal can be the same vector but pointing 180 degrees in the other direction

OpenStudy (chrisplusian):

@ybarrap I am not getting zero as the dot product, to be sure I am doing it correctly..... when I dot a vector with another vector I add the corresponding elements correct?

OpenStudy (ybarrap):

So try the negative of what you have

OpenStudy (theeric):

I don't remember that... But they are both normal, and it's worth a shot if it might be correct.

OpenStudy (chrisplusian):

@ybarrap how would you check to be sure the normal is pointing in the correct direction?

OpenStudy (chrisplusian):

I know I can use the right hand rule for a cross product but how do you know for a dot product?

OpenStudy (chrisplusian):

Call me old fashioned.... I just took \[<0,1,0> <-\frac{ 1 }{ \sqrt{2} },0,\frac{ 1 }{ \sqrt{2} }>\]

OpenStudy (chrisplusian):

I couldn't find the dot operator

OpenStudy (theeric):

I type them, and I just use `\bullet`, like \(\vec A\bullet\vec B\)

OpenStudy (ybarrap):

Where did you get <0,1,0> is that the tangent at (7,0,0)?

OpenStudy (chrisplusian):

That is the unit tangent vector evaluated at t=0

OpenStudy (theeric):

Also, I use `\left< stuff \right>` for the \(\left< stuff \right>\)

OpenStudy (ybarrap):

\cdot works also: \(\Huge \cdot \)

OpenStudy (ybarrap):

So you get zero, we agree. I would try negative.

OpenStudy (theeric):

\(\vec A\cdot\vec B\) Awesome, thanks! :)

OpenStudy (chrisplusian):

I don't see how I got zero because by my calulations I did not

OpenStudy (theeric):

Just look at where the zeroes are, actually!

OpenStudy (theeric):

\(\left<0,a,0\right>\cdot\left<b,0,d\right>\) Doesn't matter what your numbers are! Each component has a \(0\).

OpenStudy (theeric):

\(\left<0,a,0\right>\cdot\left<b,0,d\right>=0b+a0+0d=0\)

OpenStudy (chrisplusian):

I see why I didn't get zero, I messed up but how would you decide if the normal is pointing in the right direction?

OpenStudy (chrisplusian):

well that was my last attempt, I tried to change the signs, but that was wrong too!!Arghhhhhh Thanks for the help you guys, I wonder if we are doing something wrong, or if it has the wrong answer programmed in

OpenStudy (theeric):

It could just be that it has the formatting done incorrectly. I'm pretty confident in our solution, mostly because there are three of us.

OpenStudy (ybarrap):

Also, it didn't ask for unit vectors, just vectors. Maybe?

OpenStudy (theeric):

So, \(\vec T'(t)\) but not \(\dfrac{\vec T'(t)}{\left|\left|\vec T'(t)\right|\right|}\)?